On January 1, 2020, Mania Enterprises issued 10-year. 12% bonds dated January 1, 2020. The bonds have a face amount of $200,000, and they pay interest semiannually on June 30 and December 31. At the date the bonds were issued, the market yield for similar bonds was 10%.

Required: 1. Create a table like the shown one below on your answer sheet. Enter the information given in this problem into that table. Enter the words, “Solve for” to indicate which of these five variables you think you are supposed to compute. N I% PV PMT Timing Note: Timing is either “beginning” or “end.” FV

2. Compute the issue price of these bonds.

3. Prepare the journal entry to record the issuance of these bonds.
Answer #1

Calculation is shown below:

A B с D E 1 2 The price of bond needs to be computed and thus variable PV will have to be solved 3 N 20 4 1% 5.00% 5 PMT $12.

Working is shown below:

B с D =10*2 =10%/2 =200000*(12/2) End 200000 =PV(B4,B3-B5,-B7,0) Solve A 1 2 The price of bond needs to be computed and thus